1.15 mol HF is added to enough 0.202-M NaF solution to give a final volume of 2.7 L. What is the pH of the resulting solution given that the Ka of HF is 3.5x10-4 under these conditions?

2.76

1.75

0.37

3.50

3.13
Answer #1

[HF] = mol of HF / volume in L

= 1.15 mol / 2.7 L

= 0.426 M

Ka = 3.5*10^-4

pKa = - log (Ka)

= - log(3.5*10^-4)

= 3.456

use:

pH = pKa + log {[conjugate base]/[acid]}

= 3.456+ log {0.202/0.426}

= 3.132

Answer: 3.13